OpenStudy (anonymous):

write as a mixed number in lowest terms: 19.8

OpenStudy (anonymous):

\[19.8=\frac{ 19.8 }{1 }=\frac{ 19.8 }{1 }(\frac{ 10 }{10 })=\frac{ 198 }{10 }\]

OpenStudy (anonymous):

\[ 19.8 = \frac{198}{10} \]They're both even though so:\[ = \frac{99}{5} \]We know \(5\) is a prime, so there's no way this is going to reduce unless \(5\) is a factor of \(99\).

OpenStudy (anonymous):

For mixed, you need to find remainder and divisor

OpenStudy (anonymous):

why wouldnt the 19 be a whole number or part of the solution?

OpenStudy (anonymous):

I mean divident

OpenStudy (anonymous):

Yeah dividend it 19 so

OpenStudy (anonymous):

\[ 19\frac{8}{10} = 19\frac{4}{5} \]

OpenStudy (anonymous):

I was just getting them started. 198/10 gets rid of the decimal, giving you\[19\frac{ 8 }{ 10 }\] reducing to \[19\frac{ 4 }{5 }\]

OpenStudy (anonymous):

oh ok i see now! how would you do 20.2 into the same thing lowest term? does the same rule apply like reducing it

OpenStudy (anonymous):

Multiply top and bottom by 10 to get rid of the decimal then reduce. Or you can see that 20.2 is the same as 20+2/10, which is 20+1/5.